How can I find the maximum area of 4 individual pig pens with a total perimeter of 500m?
It looks something like this
_________
|____|____|
|____|____|x
y

Answer by khwang(438)

We see that the perimeter is 6x+6y = 500.
And the area is A = 4xy.

From 6x+6y = 500, we have y = 250/3 - x.
So, A = 4x(250/3 - x) = -4[x^2 -250 x/3 + (125/3)^2] + 4*125^2/9
= -4(x -125/3)^2 + 62500/9
Since -4(x -125/3)^2 < =0,
we see that when x = 125/3 the area A has max value 62500/9.
That is , when x = y = 125/3,A has max 4xy = 62500/9.
Kenny